What is the weight of a steel plate 15 feet long, 4 feet wide, and 5/8 inches

Thick, if the steel weighs 475 lbs/ ft3 .

  1. Find the Lateral surface are of the steel plate in square feet.
  2. Compute the Total surface area in square feet.





1
Expert's answer
2022-04-16T04:13:40-0400

For the first convert all units to feet: 5/8 inches is 0.05208 feet.


Volume of the plate is V = length*width*thickness = 15*4*0.05208 = 3.1248 ft3.


Mass of the plate is m = V*density = 3.1248*475 = 14 84.28 lbs.


Lateral surface are of the steel plate is

2*length*thickness+2*width*thickness = 2*15*0.05208+2*4*0.05208 = 1.979 ft2.


Total surface area of the plate is

2*length*width+2*length*thickness+2*width*thickness = 2*15*4+2*15*0.05208+2*4*0.5208 = 121.979 ft2.